A Bibliography of Publications in Notices of the American Mathematical Society (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 August 2024 Version 1.00

Title word cross-reference

#6688 [Alt84].

L [Zag84].

-series [Zag84].

1983 [Ano83]. 1984 [Ano84]. 1985 [Ano85b, Ano85a]. 1986 [AB86, Ano86c]. 1987 [Ano87]. 1988 [Ano88b, Ano88a]. 1989 [Ano89b, Ano89a].

Addendum [AV90]. advances [Rum83]. AIDS [Jac89]. Albany [Ano83]. algebraic [Qui86]. algebras [Jon86]. Algorithms [Zal88]. along [Alt84]. Amer. [Alt84].

AMS [Ano88c, LeV88]. Appendix [Jaf84]. applications [Tau86]. April [AB86]. arithmetic [Cox85]. Author [Doo89]. Award [Ano88d]. awarded [Ano83, Ano84, Ano85b, Ano85a, Ano86c, Ano87, Ano88b, Ano88a, Ano89b, Ano89a]. axis [Hal83].

Bôcher [Ano89b]. Barkley [Sac89]. barrier [Rot89]. BDF [Hal83]. been [GW89]. Bieberbach [Fit85]. Bing [AB86]. Birch [Zag84]. Birkhoff [Ano88b]. Boulder [Ano89a]. bounds [Hås88]. Bourbaki [Die82]. Bowman [HLL^{+}85].

C [Jaf84]. celebration [Ano88a]. centennial [Ano88a]. Chaitin [Del89]. circle [GW89]. City [Ano87]. class [Zag84].

Cole [Ano85b]. Combinatorics [Bjö85]. Compatibles [BP87, otB86]. complexity [Hås88]. computational [Hås88]. conjecture [Fit85, Har86, Zag84]. Constructions [Bie89]. critical [Jaf84]. curves [Zag84].

differential [Alt84, Str82]. dimensional [Fre84]. Director [Ano88e]. Document [Lam87]. does [Kis89]. during [Die82]. Dyer [Zag84]. dynamical [Gri86].

Edward [FK89]. Elliott [Ano88b]. elliptic [Zag84]. English [Alt84]. epidemic [Jac89]. equation [Del89]. equations [Alt84]. Erratum [Alt84]. Eugene [Ano84]. Everett [Ano88c]. Executive [Ano88e]. extension [Del89].

Factoring [Lan87]. first [Ano88d]. forward [Str82]. four [Fre84]. four-dimensional [Fre84]. Fourier [Str82]. Freedman [Ano86c]. Friedman [NH84]. Friend [Slo86]. future [Jaf84].

gauge [Tau86]. Gauss [Cox85, Zag84]. geometric [Cox85]. George [Ano85b]. Gödel [Del89]. Gordon [LeV89]. group [Bie89].

H [AB86, Ano86c, Ano88b]. happen [Kis89]. Harvey [Mac89, NH84].

IBM [otB86, BP87]. II [Ros82]. IIB [BP87]. infinite [Hal83]. insights [Jac89]. introduction [Pal86a].

J [Ano88e]. James [FK89]. John [Sac89]. Julia [HLL^{+}85].

Kantorovich [Ano86a]. knot [Jon86]. Krasnosel’skii [Alt84].

L. [Ano86a]. laboratory [Ash86]. Lake [Ano87]. Langlands [Ano88d]. Laramie [Ano85a]. last [Die82, GW89]. later [AV89, AV90]. LeVeque [Ano88e]. Lieb [Ano88b]. Local [Qui86]. Loftis [LeV89]. Logical [Lam87]. long [GW89]. Lower [Hås88]. Lusztig [Ano85b].

M [Ano89b]. M. [Alt84]. MACSYMA [Slo86]. Manifesto [Sie86]. Marshall [Mac89]. Math [Alt84]. Mathematica [Her88]. Mathematical [Gri86, Pal86b, Jac89, Tau86]. mathematicians [Ros82]. Mathematics [Ano86b, Ros82, Jaf84, Ano88d]. matrix [Ash86]. McShane [FK89]. mean [Cox85]. meaning [Rot89]. meeting [Ano84, Ano85a, Ano87, Ano89a]. memorium [Rot89]. method [Str82]. Michael [Ano86c]. Modeling [Jac89]. Mordell [Har86]. MR [Kis89]. MR0223640 [Alt84]. My [Slo86].

NAS [Ano88d]. networks [Bie89]. Neumann [Jon86]. Newton [AV89, AV90]. no [Fre84]. number [Zag84].

October [AB86]. operator [Alt84]. operators [Str82]. Ordering [Jaf84]. Oscar [Mum86]. other [Bot85].

P [Ano88d]. paper [Kis89]. Para [Str82]. Para-differential [Str82]. Part [BP87]. Patentable [Zal88]. PC [otB86, SW88]. Physical [Tau86]. Pitcher [Ano88c]. polynomial [Jon86]. polynomials [Lan87]. Pre [Sie86]. Pre-TEX [Sie86]. primality [Rum83]. principal [Hal83]. Principia [AV89, AV90]. prize [Ano85b, Ano89b, Ano86c, Ano88b, Knu87]. prizes [Ano83, Ano87, Ano84, Ano85a, Ano88a, Ano89a]. problem [Zag84]. Processing [Pal86b]. Production [Lam87]. provide [Jac89]. Providence [Alt84, Ano88a]. published [Kis89].

Publisher [Gil89].

qualitative [Jac89]. quickly [Lan87].

R. [AB86]. R.I. [Alt84]. read [AV89, AV90]. receives [Ano88d]. representations [Bie89]. resource [Gri86, Jaf84]. Response [Knu87]. retires [Ano88c, Ano88e]. retrospective [Fit85]. Review [Her88, Kis89]. Reviewing [Jaf84]. Reviews [BP87]. Richard [Ano89b]. Robert [Ano88d]. Robinson [HLL^{+}85]. role [Jaf84]. room [Fre84]. Rosser [Sac89].

Salt [Ano87]. Schoen [Ano89b]. science [Ano86b]. sciences [Gri86]. Secretary [Ano88c]. series [Zag84]. Soc [Alt84]. soon [LeV88]. space [Fre84]. spare [Fre84]. squared [GW89]. Stanisław [Rot89]. Steele [Ano84, Ano85a, Ano88a, Ano89a, Ano83, Ano87, Knu87]. step [Str82]. Stone [Mac89]. summer [Ano84, Ano85a, Ano87, Ano89a]. Supercalculators [SW88]. Swinnerton [Zag84]. Swinnerton-Dyer [Zag84].

Technical [otB86]. techniques [Jac89]. telephone [Bie89]. testing [Rum83]. Text [Pal86b]. theorem [Del89, Hal83]. theories [Tau86]. theory [Hås88]. There [Fre84]. things [Bot85]. thirty [Die82]. thread [Ano86b]. topology [Bjö85, Bot85, Qui86]. trajectories [Alt84]. translation [Alt84]. Typesetting [Doo89, Gil89, Sie86].

U.S. [Jaf84]. Ulam [Rot89]. unifying [Ano86b, Gri86]. universe [Jaf84]. Use [Sie86].

V [Ano86a]. Veblen [Ano86c]. View [Doo89, Gil89]. Visual [Lam87].

Walker [LeV89]. War [Ros82]. Wider [Sie86]. William [Ano88e].

Wordprocessors [BP87, otB86]. work [Die82, NH84]. World [Ros82].

years [AV89, AV90, Die82].

Zariski [Mum86].

References

Anderson:1986:RHB

[AB86] R. D. Anderson and C. E. Burgess. R. H. Bing: October 20, 1914– April 28, 1986. Notices of the American Mathematical Society, 33 (4):595–596, August 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Altshuller:1984:EOT

[Alt84] Dmitry A. Altshuller. Erratum: The operator of translation along the trajectories of differential equations [English translation, Amer. Math. Soc., Providence, R.I., 1968; MR0223640 (36 #6688)] by M. A. Krasnosel’skiĭ. Notices of the American Mathematical Society, 31 (6):621–622, October 1984. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1983:SPA

[Ano83] Anonymous. 1983 Steele prizes awarded in Albany. Notices of the American Mathematical Society, 30 (6):574–579, October 1983. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1984:SPA

[Ano84] Anonymous. 1984 Steele Prizes awarded at summer meeting in Eugene. Notices of the American Mathematical Society, 31(6):567– 569, October 1984. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1985:SPA

[Ano85a] Anonymous. 1985 Steele Prizes awarded at summer meeting in Laramie. Notices of the American Mathematical Society, 32(5):575– 579, October 1985. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

https://www.ams.org/journals/
notices/198510.

Anonymous:1985:GLA

[Ano85b] Anonymous. George Lusztig awarded 1985 Cole prize. Notices of the American Mathematical Society, 32(2):159, March 1985. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1986:VK

[Ano86a] Anonymous. L. V. Kantorovich (1912–1986). Notices of the American Mathematical Society, 33(3): 503–504, June 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1986:MUT

[Ano86b] Anonymous. Mathematics: the unifying thread in science. Notices of the American Mathematical Society, 33(5):716–733, October 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). A symposium with contributions by Phillip A. Griffiths, Allan M. Cormack, Herbert A.

Hauptman, Steven Weinberg and Isadore M. Singer.

Anonymous:1986:MHF

[Ano86c] Anonymous. Michael H. Freedman awarded 1986 Veblen Prize. Notices of the American Mathematical Society, 33(2):227–228, March 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1987:SPA

[Ano87] Anonymous. 1987 Steele prizes awarded at the summer meeting in Salt Lake City. Notices of the American Mathematical Society, 34 (6):875–878, October 1987. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1988:SPA

[Ano88a] Anonymous. 1988 Steele Prizes awarded at centennial celebration in Providence. Notices of the American Mathematical Society, 35(7): 965–970, September 1988. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1988:EHL

[Ano88b] Anonymous. Elliott H. Lieb awarded 1988 Birkhoff Prize. Notices of the American Mathematical Society, 35(3):381–382, March 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1988:EPR

[Ano88c] Anonymous. Everett Pitcher retires as AMS Secretary. Notices of the American Mathematical Society, 35(10):1463–1466, December 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1988:RPL

[Ano88d] Anonymous. Robert P. Langlands receives first NAS Award in Mathematics. Notices of the American Mathematical Society, 35(4): 509–510, April 1988. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1988:WJL

[Ano88e] Anonymous. William J. LeVeque retires as Executive Director. Notices of the American Mathematical Society, 35(6):783–784, July/ August 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). With comments by George Daniel Mostow and Irving Kaplansky.

Anonymous:1989:SPA

[Ano89a] Anonymous. 1989 Steele Prizes awarded at Summer Meeting in Boulder. Notices of the American Mathematical Society, 36(7): 831–836, September 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Anonymous:1989:RMS

[Ano89b] Anonymous. Richard M. Schoen awarded 1989 Bôcher prize. Notices of the American Mathematical Society, 36(3):225–226, March 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Ash:1986:ML

[Ash86] A. Ash. A matrix laboratory. Notices of the American Mathematical Society, 33(3):482–485, June 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Arnold:1989:NPR

[AV89] V. I. Arnol’d and V. A. Vasil’ev. Newton’s Principia read 300 years later. Notices of the American Mathematical Society, 36(9): 1148–1154, November 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See addendum [AV90].

Arnold:1990:ANP

[AV90] V. I. Arnol’d and V. A. Vasil’ev. Addendum to: “Newton’s Principia read 300 years later”. Notices of the American Mathematical Society, 37(2):144, February 1990. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). See [AV89].

Barwise:1989:TLA

[Bar89] Jon Barwise. TEX, LaTEX, and AMS-TEX. Notices of the American Mathematical Society, 36 (3):241–243, March 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bien:1989:CTN

[Bie89] Frédéric Bien. Constructions of telephone networks by group representations. Notices of the American Mathematical Society, 36(1):5– 22, January 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bjorner:1985:CT

[Bjö85] Anders Björner. Combinatorics and topology. Notices of the American Mathematical Society, 32(3): 339–345, June 1985. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Bott:1985:TOT

[Bot85] Raoul Bott. On topology and other things. Notices of the American Mathematical Society, 32(2):152– 158, March 1985. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

BCS:1987:TWI

[BP87] Boston Computer Society and PC Technical Group of IBM PC Users Group. Technical wordprocessors for the IBM PC and compatibles, Part IIB: Reviews. Notices of the American Mathematical Society, 34(3):462–491, April 1987. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Cox:1985:GAG

[Cox85] David A. Cox. Gauss and the arithmetic–geometric mean. Notices of the American Mathematical Society, 32(2):147–151, May 1985. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Delahaye:1989:CEE

[Del89] Jean-Paul Delahaye. Chaitin’s equation: an extension of Gödel’s theorem. Notices of the American Mathematical Society, 36(8):984– 987, October 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Dieudonne:1982:WBD

[Die82] Jean Dieudonné. The work of Bourbaki during the last thirty years. Notices of the American Mathematical Society, 29(7):618–623, November 1982. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Doob:1989:TTA

[Doo89] Michael Doob. TEX and typesetting — an author’s view. Notices of the American Mathematical Society, 36(9):1203–1204, November 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

FitzGerald:1985:BCR

[Fit85] Carl H. FitzGerald. The Bieberbach conjecture: retrospective. Notices of the American Mathematical Society, 32(1):2–6, January 1985. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fleming:1989:EJM

[FK89] Wendell Fleming and Victor Klee. Edward James McShane (1904– 1989). Notices of the American Mathematical Society, 36(7): 828–830, September 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Freedman:1984:TNR

[Fre84] Michael H. Freedman. There is no room to spare in four-dimensional space. Notices of the American Mathematical Society, 31(1):3– 6, January 1984. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gilchrist:1989:TTP

[Gil89] Martin Gilchrist. TEX and typesetting — a publisher’s view. Notices of the American Mathematical Society, 36(9):1199–1201, November 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Griffiths:1986:MSU

[Gri86] Phillip A. Griffiths. Mathematical sciences: a unifying and dynamical resource. Notices of the American Mathematical Society, 33(3): 463–479, June 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Gardner:1989:LLC

[GW89] Richard J. Gardner and Stan Wagon. At long last, the circle has been squared. Notices of the American Mathematical Society, 36(10): 1338–1343, December 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Halmos:1983:BIP

[Hal83] P. R. Halmos. BDF or the infinite principal axis theorem. Notices of the American Mathematical Society, 30(4):387–391, June 1983. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Harris:1986:MC

[Har86] David Harris. The Mordell conjecture. Notices of the American Mathematical Society, 33(3): 443–449, June 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Haastad:1988:LBC

[Hås88] Johan Håstad. Lower bounds in computational complexity theory. Notices of the American Mathematical Society, 35(5):677–683, May/ June 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Herman:1988:MR

[Her88] Eugene A. Herman. Mathematica — a review. Notices of the American Mathematical Society, 35(9): 1334–1344, November 1988. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Henkin:1985:JBR

[HLL^{+}85] L. Henkin, D. H. Lehmer, Emma Lehmer, E. Scott, J. Kelley, L. Gaal, D. Gale, M. Davis, S. MacLane, I. Niven, E. Pitcher, L. Blum, and S. Feferman. Julia Bowman Robinson: 1919–1985. Notices of the American Mathematical Society, 32(6):739–742, November 1985. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jackson:1989:MAE

[Jac89] Allyn Jackson. Modeling the AIDS epidemic: mathematical techniques provide qualitative insights. Notices of the American Mathematical Society, 36(8):981–983, October 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jaffe:1984:RUM

[Jaf84] Arthur M. Jaffe. Reviewing U.S.

mathematics: critical resource for the future. Appendix C. Ordering the universe: the role of mathematics. Notices of the American Mathematical Society, 31(6):589–608, October 1984. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Jones:1986:NKP

[Jon86] V. F. R. Jones. A new knot polynomial and von Neumann algebras. Notices of the American Mathematical Society, 33(2):219–225, March 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Kister:1989:PPM

[Kis89] Jane E. Kister. From published paper to MR review: how does it happen? Notices of the American Mathematical Society, 36(10): 1362–1365, December 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Knuth:1987:RSP

[Knu87] Donald E. Knuth. Response to the Steele Prize. Notices of the American Mathematical Society, 34 (2):227–228, March 1987. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Lamport:1987:DPV

[Lam87] Leslie Lamport. Document production: Visual or logical? Notices of the American Mathematical Society, 34(4):621–624, June 1987. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Landau:1987:FPQ

[Lan87] Susan Landau. Factoring polynomials quickly. Notices of the American Mathematical Society, 34(1): 3–8, January 1987. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LeVeque:1988:ATN

[LeV88] William J. LeVeque. The AMS — then, now, and soon. Notices of the American Mathematical Society, 35(6):785–789, July/August 1988. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

LeVeque:1989:GLW

[LeV89] William J. LeVeque. Gordon Loftis Walker 1912–1988. Notices of the American Mathematical Society, 36(2):139, February 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mackey:1989:MHS

[Mac89] George W. Mackey. Marshall Harvey Stone. 1903–1989. Notices of the American Mathematical Society, 36(3):221–223, March 1989. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Mumford:1986:OZ

[Mum86] David Mumford. Oscar Zariski:

1899–1986. Notices of the American Mathematical Society, 33(6): 891–894, November 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Nerode:1984:WHF

[NH84] Anil Nerode and Leo A. Harrington. The work of Harvey Friedman. Notices of the American Mathematical Society, 31(6):563–566, October 1984. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

BCS:1986:TWI

[otB86] Report of the Boston Computer Society. Technical Wordprocessors for the IBM PC and Compatibles. Notices of the American Mathematical Society, 33(1):8–37, January 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Palais:1986:ITA

[Pal86a] Richard S. Palais. An introduction to TEX and AMS-TEX. Notices of the American Mathematical Society, 33(2):303–308, March 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Palais:1986:MTP

[Pal86b] Richard S. Palais. Mathematical text processing. Notices of the American Mathematical Society, 33(1):3–7, January 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Quinn:1986:LAT

[Qui86] Frank Quinn. Local algebraic topology. Notices of the American Mathematical Society, 33(6): 895–899, November 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rosser:1982:MMW

[Ros82] J. Barkley Rosser. Mathematics and mathematicians in World War II. Notices of the American Mathematical Society, 29(6):509–515, October 1982. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rota:1989:BMM

[Rot89] Gian-Carlo Rota. The barrier of meaning. In memorium: Stanisław Ulam. Notices of the American Mathematical Society, 36(2):141– 143, February 1989. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rumely:1983:RAP

[Rum83] Robert Rumely. Recent advances in primality testing. Notices of the American Mathematical Society, 30 (5):475–477, August 1983. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sacks:1989:JBR

[Sac89] Gerald E. Sacks. John Barkley Rosser (1907–1989). Notices of the American Mathematical Society, 36 (10):1367, December 1989. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Siebenmann:1986:TWU

[Sie86] L. C. Siebenmann. Toward wider use of TEX typesetting. A pre-TEX manifesto. Notices of the American Mathematical Society, 33(4):597– 607, August 1986. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Sloane:1986:MFM

[Slo86] N. J. A. Sloane. My friend MAC-SYMA. Notices of the American Mathematical Society, 33(1):40–43, January 1986. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Strichartz:1982:DOA

[Str82] Robert S. Strichartz. Paradifferential operators — another step forward for the method of Fourier. Notices of the American Mathematical Society, 29(5):402– 406, August 1982. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Simon:1988:SP

[SW88] Barry Simon and Richard M. Wilson. Supercalculators on the PC. Notices of the American Mathematical Society, 35(7):978–1001, September 1988. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taubes:1986:PMA

[Tau86] Clifford Henry Taubes. Physical and mathematical applications of gauge theories. Notices of the American Mathematical Society, 33 (5):707–715, October 1986. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zagier:1984:SEC

[Zag84] D. Zagier. L-series of elliptic curves, the Birch–Swinnerton-Dyer conjecture, and the class number problem of Gauss. Notices of the American Mathematical Society, 31 (7):739–743, November 1984. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Zalta:1988:AP

[Zal88] Edward N. Zalta. Are algorithms patentable? Notices of the American Mathematical Society, 35(6): 796–799, July/August 1988. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).